A solution contain 25% ethanol, 50% acetic acid and 25% water by mass. Calculate the mole fraction of each components ????

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer :

For simplicity take: 25g H2O . 25g ethanol and 50g CH3COOH
Calculate mol of each:
molar mass H2O = 18g/mol
25g = 25/18 = 1.39 mol
Molar mass ethanol CH3CH2OH = 46g/mol
25g = 25/46 = 0.543 mol
Molar mass CH3COOH = 60g/mol
50g = 50/60 = 0.833 mol
Total mol = 1.39+0.543 +0.833 = 2.766 mol
Mol fraction H2O = 1.39/2.766 = 0.503
Mol fraction CH3CH2OH = 0.196
Mol fraction CH3COOH = 0.833/2.766 0.301

Thanks